Jasmine purchased a machine for $29,500 for her company. She paid 5.00% of this amount as a down payment and financed the rest at 6.12% compounded monthly. She paid $1,125 at the end of every month to settle the loan.

What was the principal portion of payment number 4?

What was the interest portion of payment number 4?
Answer #1

Loan 29500*95% $ 28,025.00 G=a*0.0612/12 Interest d=b-c Pricipal Month Beginning Loan Balance $28,025.00 $27,042.93 $26,055.8